myininaya (myininaya):

provide proof or counterexample here: if x and y have no common factors, then x+y is prime

myininaya (myininaya):

x and y are integers

OpenStudy (anonymous):

x = 3 y = 7 x + y = 10, not prime

myininaya (myininaya):

ok lol

myininaya (myininaya):

lets also suppose x and y are not prime

OpenStudy (anonymous):

if x and y had a common factor, say r, then x = rt, y = rq x + y = r(t + q) which isnt prime.

OpenStudy (anonymous):

all variables are integers btw

myininaya (myininaya):

but x and y share r

myininaya (myininaya):

wait are you trying to do a contradiction

OpenStudy (anonymous):

if x and y aren't prime... x = 4 y = 45 so x and y have no common factors, and they're both composite x + y = 49, which isn't prime

OpenStudy (anonymous):

...i think i am lol

myininaya (myininaya):

nice work you guys im gonna come up with something harder

myininaya (myininaya):

you'll see

myininaya (myininaya):

i just made an evil laugh

OpenStudy (anonymous):

lol

myininaya (myininaya):

joe you still there i got one for you i think you might like its not a proof kindof question

OpenStudy (anonymous):

sure thing, fire away

myininaya (myininaya):

Let n be an integer such that if n is divided by 3, then the remainder is 2. and if n is divided by 5, then the remainder is 1. (There can be more than one n, but give the smallest value of n that satisfies the above)

myininaya (myininaya):

n is a positive integer

OpenStudy (anonymous):

i got an answer....since i sorta kinda know about these type of problems (Chinese Remainder Theorem I believe), but I dont know the correct way to do them >.< if i turned in my solution for a test or homework, im pretty sure it would get counted wrong.

OpenStudy (anonymous):

gonna post my solution, one sec

myininaya (myininaya):

thats right to use chinese remainder thm

OpenStudy (anonymous):

26? just a guess

myininaya (myininaya):

11

OpenStudy (anonymous):

that works, not the smallest though. really nice guess btw :)

myininaya (myininaya):

yes nice guess

OpenStudy (anonymous):

its any number congruent to 11 mod 15

OpenStudy (anonymous):

ah yes, nice joe. n = 3k + 2 n = 5m + 1 so 3k + 2 = 5m + 1 3k + 1 = 5m the smallest solution of which is k = 3, m = 2

OpenStudy (anonymous):

alright, here's my solution, beware, it butchers a lot of properties >.< my excuse is that i haven't learned this in a class yet, i was reading about it and i sotra kinda understand whats going on, but not fully >.>

OpenStudy (anonymous):

the 15 comes from 3*5, i should have put that somewhere on there

myininaya (myininaya):

\[\frac{n}{3}=Q_1+\frac{2}{3}\] \[\frac{n}{5}=Q_2+\frac{1}{5}\] so we have \[n=3Q_1+2\] and \[n=5Q_2+1\] \[3Q_1+2=5Q_2+1\] \[5Q_2-3Q_1=1\] ---- \[5=3*1+2\] \[3=2*1+1\] ---- so 3-2=1 3-(5-3)=1 -1*5+2*3=1 -1*5+(5-3)*3=1 2*5-3*3=1 so \[Q_2=2, Q_1=3\] but we only need one of these to find n --- so \[n=5Q_2+1=5*2+1=10+1=11\]

myininaya (myininaya):

i use euclidean

myininaya (myininaya):

then i went backwards

OpenStudy (anonymous):

Ah, that is a much more elegant solution, is that Bezout's Identity in there?

myininaya (myininaya):

im not sure the only thing i know for sure that i used was euclidean i dont know of bezout's identity

myininaya (myininaya):

im not sure what the chinese thm says but i think it has something to do with euclidean or gcd(5,3)=1 then we have 5x+3y=1 i think

OpenStudy (anonymous):

its that if gcd(a, b) = 1, then there are some integers x and y such that: ax + by = 1 it sorta looks like your 5Q_2-3Q_1 = 1

myininaya (myininaya):

right i do remember it :)

OpenStudy (anonymous):

er, rather, thats a special case, its really if the gcd (a,b) = d, then there are some integers x and y such that: ax+by = d

myininaya (myininaya):

i took 5 sense it was bigger than 3 and did 5/3 5=3*2+1 3=2*1+1 and after this i went backwards 3-2=1 3-(5-3)=1 and blah blah

OpenStudy (anonymous):

nice nice, i like that way better than the way i learned it.

myininaya (myininaya):

oops 5=3(1)+2 3=2(1)+1

myininaya (myininaya):

so you are taking discrete?

myininaya (myininaya):

no this was number theory lol

myininaya (myininaya):

number theory*

OpenStudy (anonymous):

ive been reading some number theory books lately, im registered to take the class this fall, but its not required to get a BS in Math, and as such, not many people at my school take it. It will be a miracle if enough people sign up for it and it doesnt get cancelled =/

myininaya (myininaya):

omg you will not be dissatisfied if you take this class i love number theory i wish i had more of it i'm trying to do research but it is so hard when someone is not making you do this stuff

myininaya (myininaya):

i almost wasted the whole summer goofing off i need to do it

OpenStudy (anonymous):

It does seem extremely interesting, in one of the books there is this section on something called Diophantine (i think i spelled that right) equations, and i was totally in love lol.

myininaya (myininaya):

lol i have to get ready for bed goodnight joe

OpenStudy (anonymous):

goodnight, see ya later